# T-SCEND: Test-time Scalable MCTS-enhanced Diffusion Model

### v2

@@ Abstract
We introduce Test-time Scalable MCTS-enhanced Diffusion Model (T-SCEND), a novel framework that significantly improves diffusion model’s reasoning capabilities with better energy-based training and scaling up test-time computation. We first show that naïvely scaling up inference budget for diffusion models yields marginal gain. To address this, the training of T-SCEND consists of a novel linear-regression negative contrastive learning objective to improve the performance-energy consistency of the energy landscape, and a KL regularization to reduce adversarial sampling. During inference, T-SCEND integrates the denoising process with a novel hybrid Monte Carlo Tree Search (hMCTS), which sequentially performs best-of-N random search and MCTS as denoising proceeds. On challenging reasoning tasks of Maze and Sudoku, we demonstrate the effectiveness of T-SCEND’s training objective and scalable inference method. In particular, trained with Maze sizes of up to $6\times 6$, our T-SCEND solves 88% of Maze problems with much larger sizes of $15\times 15$, while standard diffusion completely fails. Code to reproduce the experiments can be found at https://github.com/AI4Science-WestlakeU/t_scend.
@@ 1 Introduction
In recent years, diffusion generative models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021) have demonstrated remarkable performance across a wide range of applications, including image (Rombach et al., 2022) and video generation (Brooks et al., 2024), robotic control (Janner et al., 2022), and reasoning (Du et al., 2024). By learning how to denoise from noisy, corrupted samples to clean data samples, diffusion models can generate high-quality, high-dimensional samples approximating the data distribution.
@@
Figure 1: Visualizations of Maze training data and solutions generated by hMCTS denoising of our T-SCEND framework.
Nevertheless, as the application scenarios for diffusion models become more complex, they often encounter situations where the problem sizes are much larger than those in training or their conditional input is out-of-distribution (OOD). When dealing with larger problem sizes, the computational resources required may increase exponentially, and the existing models may struggle to maintain high-quality performance. In OOD scenarios, the diffusion models need to generalize to conditions that have characteristics different from what they have been trained on. Thus, the pressing question is how to further enhance the performance of diffusion models under such challenging circumstances.
Interestingly, test-time scaling has demonstrated its effectiveness in other generative AI methods such as large language models (Wei et al., 2022; Yao et al., 2024). It remains an interesting question whether test-time [^1] In this paper, we use “test-time” and “inference” interchangeably. scaling can also benefit diffusion models and address the challenges above.
However, naïve test-time scaling up for diffusion models by increasing denoising steps typically yields marginal gain after a certain number of function evaluations (NFEs), which was observed in prior works (Karras et al., 2022; Song et al., ; Song et al., 2021). Another scaling-up dimension is the number of samples, while in our pivot study, best-of-N sampling with diffusion model still doesn’t scale up significantly. Through analysis, we identify that the reason comes from both training and inference. For training, the model’s performance is fundamentally upper-limited by the learned energy landscape [^2] We can regard diffusion model’s denoising network $\bm{\epsilon}_{\theta}(\bm{x}_{t},t)$ as approximating the gradient of an implicit energy function $\nabla_{\bm{x}_{t}}E_{\theta}(\bm{x}_{t},t)$. The energy function connects to the probability distribution via $p(\bm{x})\propto e^{-E_{\theta}(\bm{x},t=0)}$. Thus, the lower the energy, the higher the probability. See Du et al. (2023) for more details.. For example, if the energy landscape has many local minima or has the wrong global minima, simply scaling up the inference budget will likely not result in improved performance. We identify that the insufficient performance-energy consistency and adversarial sampling [^3] Detailed description for performance-energy consistency and adversarial sampling can be found in Appendix A. as two key reasons limiting the performance. These factors make the sampling easier to be stuck at sub-optimal solutions and hard to go out. For inference, simple search methods such as random search may not be efficient enough to navigate the vast space when generating high-dimensional samples.
In this work, we introduce Test-time Scalable MCTS-enhanced Diffusion Model (T-SCEND), a novel framework that significantly improves diffusion model’s reasoning capability. T-SCEND consists of innovations in both training objectives and inference scaling. On the training side, T-SCEND incorporates a novel Linear-Regression Negative Contrastive Learning (LRNCL) objective to improve the performance-energy consistency of the energy landscape, and a KL regularization to address adversarial sampling. Specifically, for the LRNCL objective, it randomly samples two negative samples from a positive sample and performs a linear regression for these three samples on their energy vs. L2 distance to the positive sample. The LRNCL objective encourages that the samples’ energy is linear w.r.t. their L2 distance to the positive sample, and the slope is positive, thus making the energy landscape reflective of their distance to the ground-truth, improving performance-energy consistency. For the KL regularization, it differentiates through the denoising process, making the energy landscape easier to sample and reducing adversarial sampling.
On the inference side, we integrate the denoising process with a novel hybrid Monte Carlo Tree Search (hMCTS). It sequentially performs best-of-N random search and MCTS as the denoising proceeds. The first few steps of best-of-N random search use $L$ independent samples that denoise from the prior (Gaussian) distribution until reaching a predefined denoising step. The best sample is chosen from the $L$ samples. Then MCTS denoising follows and completes the denoising process. The random search in the initial denoising steps allows the model to reach cleaner samples with a reasonable signal-to-noise ratio, while the later MCTS denoising allows better navigation in the search space. Taken together, the hybrid MCTS denoising allows the model to significantly improve inference performance by scaling up the inference computational budget.
@@
We demonstrate the effectiveness of our T-SCEND’s training objective and scalable inference method on challenging reasoning tasks of Maze and Sudoku. In Sudoku tasks, our T-SCEND solves $43\%$ problems when conditioned on much less given digits (OOD condition), while the standard diffusion model only solves $30\%$ problems. In Maze tasks, trained with Maze sizes of up to $6\times 6$, our T-SCEND solves 88% of Maze problems with much larger sizes of $15\times 15$, while standard diffusion completely fails.
@@
In summary, our contributions are as follows: (1) We introduce Test-time Scalable MCTS-enhanced Diffusion Model (T-SCEND) as a novel framework that can scale up test-time computation for better reasoning capability. (2) We introduce Linear-Regression Negative Contrastive Learning (LRNCL) and KL regularization into the training objective to improve the energy landscape. (3) We integrate a hybrid Monte Carlo Tree Search into the denoising process, enabling test-time scaling for better reasoning accuracy. (4) We demonstrate the effectiveness of our method on challenging Sudoku and Maze datasets where the conditions are OOD or the problem sizes are much larger.
@@
Figure 2: Overview of T-SCEND. This figure illustrates the key aspects of T-SCEND by contrasting its training and inference strategies with those of the previous method. Specifically, T-SCEND introduces ${\mathcal{L}}_{\text{LRNCL}}$ and ${\mathcal{L}}_{\text{KL}}$ during training and incorporates Monte Carlo Tree Search (MCTS) during inference to enable efficient scaling.
@@ 2 Related Work
Controllable Generation of Diffusion Models: Controllable generation in diffusion models has been a significant area of research aiming to guide the generation process to produce samples that meet specific criteria or constraints. Previous work focuses on controlling the generation process through an external or internal gradient (Dhariwal & Nichol, 2021; Ho & Salimans, 2021), which requires additional training according to each new objective. T-SCEND does not rely on external guidance/classifiers or conditional training but instead utilizes sampling, simulation, and selection during inference to navigate the generation process.
Test-time scaling and MCTS: Test-time scaling involves augmenting computational resources during the inference phase to improve model performance. This approach has gained significant attention in the field of large language models (LLMs) (Brown et al., 2024; Snell et al., 2024; Wu et al., 2024a), substantially enhancing their capabilities. Techniques such as CoT (Wei et al., 2022) and TOT (Yao et al., 2024) are commonly used to incorporate explicit reasoning steps to strengthen the model’s reasoning abilities. Monte Carlo Tree Search (MCTS) (Coulom, 2006; Kocsis & Szepesvári, 2006; Gelly et al., 2006) is a heuristic search algorithm designed for decision-making processes. Recently, MCTS has been increasingly integrated with deep learning, particularly reinforcement learning, providing an efficient means of enhancing computational search during inference (Silver et al., 2016, 2017, 2018; Schrittwieser et al., 2020).
The idea of test-time scaling for diffusion models was concurrently explored by Ma et al. (2025). Their approach investigates how test-time computation budgets affect scaling performance, as well as the influence of different verifiers and scaling search algorithms. In contrast, we focus on why naïvely trained diffusion models fail and propose KL loss and LRNCL loss to enhance the diffusion model’s capacity, facilitating better scalability. For test-time scaling methods, we introduce hMCTS denoising, which innovatively integrates MCTS and the denoising process to fully unlock the scalability of the diffusion model.
@@ 3 Preliminary
The Denoising Diffusion Probabilistic Model (DDPM) (Ho et al., 2020) contains a predefined forward process to corrupt data into Gaussian noises, and a learnable reverse process to generate new samples from them. The forward process follows a Gaussian transition kernel $q(\bm{x}_{t}|\bm{x}_{t-1})=\mathcal{N}(\bm{x}_{t}|\sqrt{\alpha_{t}}\bm{x}_{t-1},(1-\alpha_{t})\bm{I}),t=1,\cdots,T$, where the noise schedule $\{\alpha_{t}\}_{t=1}^{T}$ and $T$ is set to make $\bm{x}_{T}$ follow approximately a standard Gaussian distribution. The reverse process can be learned to predict the noise from the corrupted data by minimizing
@@
where the expectation is w.r.t. $\bm{x}_{0}\sim p(\bm{x})$, $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$, and $t\sim\{1,...T\}$. $\bar{\alpha}_{t}:=\prod_{i=1}^{t}\alpha_{i}$. Eq. 1 is equivalent to optimizing a reweighted variational bound on negative log-likelihood. Without loss of generality, in this work, we parameterize $\bm{\epsilon}_{\theta}$ as the gradient of the energy function $\nabla_{\bm{x}_{t}}E_{\theta}(\bm{x}_{t},t)$ as in Du et al. (2023) [^4] This parameterization has the benefit that it simultaneously learns ${\bm{\epsilon}}_{\theta}$ as gradient of the energy $\nabla_{\bm{x}_{t}}E_{\theta}(\bm{x}_{t},t)$ and learns the energy itself as an unnormalized log-probability due to $p({\bm{x}}_{t})\propto e^{-E_{\theta}({\bm{x}}_{t},t)}$, allowing inference to utilize this “energy” as guidance.. To improve the energy landscape, Du et al. (2024) introduced a contrastive loss
@@
The reverse process starts with $\bm{x}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$, and iteratively applies the learned denoising net $\bm{\epsilon}_{\theta}$, where Song et al. introduces an adjustable noise scale $\sigma_{t}$:
$\begin{split}&\bm{x}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\frac{\bm{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{\theta}\left(\bm{x}_{t},t\right)}{\sqrt{\bar{\alpha}_{t}}}\\ &+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\bm{\epsilon}_{\theta}\left(\bm{x}_{t},t\right)+\sigma_{t}\bm{\epsilon}_{t},\,\,\bm{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})\end{split}$ (3)
Importantly, Song et al. highlights that a diffusion model trained with $\{\alpha_{t}\}_{t=1}^{T}$ can be used to sample with $\{\alpha_{\tau_{s}}\}_{s=1}^{S}$, where $\tau$ is any increasing sub-sequence of $1,2,\cdots,T$ of length $S$, which significantly accelerates sampling, and will serve as an efficient approach for MCTS simulation in Section 4.3.
@@ 4.1 Problem setup
Consider an optimization problem $\underset{\bm{x}\in{\mathcal{C}}}{\min}\mathcal{J}(\bm{x})$ with objective $\mathcal{J}:\mathbb{R}^{n}\to\mathbb{R}$ and feasible set ${\mathcal{C}}\subseteq\mathbb{R}^{n}$. We focus on cases where $\mathbb{R}^{n}$ is a high-dimensional space and a dataset $\{\bm{x}_{i}\}$ drawn from $p(\bm{x})$ supported on ${\mathcal{C}}$ is given. The dataset has two-fold effects here. First, in real-world applications, the explicit expression of ${\mathcal{C}}$ is often unavailable, and can only be learned implicitly from samples. Second, in high-dimensional problems, data prior could make the search much more efficient, as we will demonstrate in Section 5.
4.2 Training of T-SCEND
@@
Although the denoising loss ${\mathcal{L}}_{\text{MSE}}$ (Eq. 1) and the contrastive loss ${\mathcal{L}}_{\text{Contrast}}$ (Eq. 2) are effective, we find that simply scaling up inference budget by having more independent random search processes yields marginal gain (Sec. 5.1). Through deeper analysis, we find that the core reason is insufficient performance-energy consistency and adversarial sampling, which we address by introducing a novel Linear-Regression Negative Contrastive Learning loss and incorporating a KL regularization, which we detail as follows.
Linear-Regression Negative Contrastive Learning. Concretely, although the contrastive loss (Eq. 2) encourages the positive examples to have global minimum energy, it does not stipulate the landscape for the negative examples. For example, it is likely that a negative example ${\bm{x}}_{t}^{--}$ that is further apart from the positive example ${\bm{x}}_{t}$ has lower energy than a negative example ${\bm{x}}_{t}^{-}$ that is nearer. Thus, the inference process can start around ${\bm{x}}_{t}^{--}$ and can hardly move out. This will result in reduced performance-energy consistency, where lower energy at step $t$ does not necessarily correspond to a more accurate predicted solution $\hat{{\bm{x}}}_{0}$, as will be shown in Sec. 5.1. This problem can not be easily remedied by having more inference budget such as more random search processes. A more fundamental way to solve this problem is to shape the energy landscape by regularizing the energy among negative examples during training. Specifically, from a positive example ${\bm{x}}_{0}\sim p({\bm{x}})$, we sample two negative examples ${\bm{x}}_{0}^{-}$ and ${\bm{x}}_{0}^{--}$ [^5] Details to generate negative samples are in Appendix A., the latter has a larger L2 distance to ${\bm{x}}_{0}$. Specifically, the distance of ${\bm{x}}_{0}$, ${\bm{x}}_{0}^{-}$ and ${\bm{x}}_{0}^{--}$ to ${\bm{x}}_{0}$ are 00, $l_{2,0}^{-}=||{\bm{x}}_{0}^{-}-{\bm{x}}_{0}||_{2}^{2}$, and $l_{2,0}^{--}=||{\bm{x}}_{0}^{--}-{\bm{x}}_{0}||_{2}^{2}$, respectively. Then we obtain their corresponding noisy examples at step $t$ via ${\bm{x}}_{t}=\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}$, ${\bm{x}}_{t}^{-}=\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}^{-}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}$, and ${\bm{x}}_{t}^{--}=\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}^{--}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}$. Their energy are $E^{+}_{t}=E_{\theta}({\bm{x}}_{t},t)$, $E^{-}_{t}=E_{\theta}({\bm{x}}_{t}^{-},t)$, and $E^{--}_{t}=E_{\theta}({\bm{x}}_{t}^{--},t)$, respectively. To encourage a landscape where the energy difference between negative samples and the positive sample becomes smaller as they approach the positive sample, we directly constrain the relationship between the energy of different negative samples and the positive sample using linear regression. Specifically, we first apply the linear regression algorithm [^6] The details of linear regression algorithm can be found in Appendix A. to fit a line through the three points $\{(0,E_{t}^{+}),(l_{2,0}^{-},E_{t}^{-}),(l_{2,0}^{--},E_{t}^{--})\}$, which is characterized by the slope $k_{t}$ and the intercept $b_{t}$. Then, we obtain the corresponding fitted points $\{(0,\hat{E}_{t}^{+}),(l_{2,0}^{-},\hat{E}_{t}^{-}),(l_{2,0}^{--},\hat{E}_{t}^{--})\}$. We then compute the Linear-Regression Negative Contrastive Learning (LRNCL) loss as follows:
${\mathcal{L}}_{\text{LRNCL}}=$ ${\mathbb{E}}_{{\bm{x}}_{0},{\bm{x}}_{0}^{-},{\bm{x}}_{0}^{--},{\bm{\epsilon}},t}\big{[}\max(0,\gamma-k_{t})\,+$ (4)
$||E_{t}^{+}\,-\,\hat{E}_{t}^{+}$ $||_{2}^{2}+||E_{t}^{-}-\hat{E}_{t}^{-}||_{2}^{2}+||E_{t}^{--}-\hat{E}_{t}^{--}||_{2}^{2}\big{]},$

### v3

# VFScale: Intrinsic Reasoning through Verifier-Free Test-time Scalable Diffusion Model

@@ Abstract
Inspired by human SYSTEM 2 thinking, LLMs excel at complex reasoning tasks via extended Chain-of-Thought. However, similar test-time scaling for diffusion models to tackle complex reasoning remains largely unexplored. From existing work, two primary challenges emerge in this setting: (i) the dependence on an external verifier indicating a notable gap from intrinsic reasoning of human intelligence without any external feedback, and (ii) the lack of an efficient search algorithm. In this paper, we introduce the Verifier-free Test-time Scalable Diffusion Model (VFScale) to achieve scalable intrinsic reasoning, which equips number-of-sample test-time scaling with the intrinsic energy function of diffusion models as the verifier. Concretely, VFScale comprises two key innovations to address the aforementioned challenges. On the training side, VFScale consists of a novel LRNCL loss and a KL regularization to improve the energy landscape, ensuring that the learned energy function itself serves as a reliable verifier. On the inference side, VFScale integrates the denoising process with a novel hybrid Monte Carlo Tree Search (hMCTS) to improve search efficiency. On challenging reasoning tasks of Maze and Sudoku, we demonstrate the effectiveness of VFScale’s training objective and scalable inference method. In particular, trained with Maze sizes of up to $6\times 6$, our VFScale solves 88% of Maze problems with much larger sizes of $15\times 15$, while standard diffusion model completely fails.
@@ 1 Introduction
Human intelligence can allocate more computation to solve harder problems, known as SYSTEM 2 thinking kahneman2011thinking . Motivated by this, large language models (LLMs) have exhibited promising performance in reasoning tasks through deliberate long Chain-of-Thought lightman2023let ; deepseekai2025deepseekr1incentivizingreasoningcapability . Alongside LLMs, diffusion models have recently emerged as a compelling alternative. By casting reasoning as an optimization problem, these models iteratively refine candidate solutions toward higher-quality outputs. This approach has demonstrated strong potential, achieving notable results on tasks such as Sudoku solving, graph connectivity, and shortest-path computation du2022learning ; du2024learning .
@@
Figure 1: Visualizations of Maze training data and solutions generated by hMCTS denoising of our VFScale framework.
Although the iterative denoising procedure of diffusion models mirrors the repeated reasoning cycles required by complex reasoning tasks, their performance declines sharply once the problem difficulty exceeds the training distribution du2022learning . Moreover, prior works show that simply increasing the number of sampling steps quickly leads to a performance plateau du2024learning . Instead, in image generation, inference-time scaling to improve performance by increasing the number of samples has proven effective ma2025inferencetimescalingdiffusionmodels . However, this strategy depends on an external verifier that is often unavailable or prohibitively expensive ouyang2022training ; lee2024rlaif . Yet human intelligence can engage in deep, reflective intrinsic reasoning without any external feedback, revealing a clear gap with such methods lombrozo2024learning ; chi1994eliciting ; huanglarge . These observations motivate our central question: Can we design a Verifier-free Test-time Scalable Diffusion Model to achieve scalable intrinsic reasoning?
In pursuit of such a test-time scalable intrinsic reasoning model, we leverage the fundamental mechanics of diffusion models. Diffusion models approximate the score function song2020score , which is equivalent to the negative gradient of the energy function. Our key insight is that the energy function, as a measure of the learned probability, can serve as an effective verifier itself. Based on the above premise, in this work, we explore equipping number-of-sample test-time scaling with diffusion models’ intrinsic energy function.
However, this strategy presents two primary challenges: (1) Firstly, using number-of-sample test-time scaling requires a verifier to evaluate and select samples. Prior work on test-time scaling has also established that verifier accuracy is crucial for performance setlur2025scalingtesttimecomputeverification ; ma2025inferencetimescalingdiffusionmodels . Consequently, using the energy function as an intrinsic verifier requires it to reliably reflect the sample quality. Nevertheless, previous energy-based diffusion models have largely ignored this requirement. (2) Secondly, the efficiency of the search algorithm is critical when scaling by varying the number of samples snell2024scalingllmtesttimecompute . Nevertheless, the search efficiency of existing methods, such as best-of-$N$ (BoN), warrants further enhancement.
To approach these challenges, we propose Verifier-Free Test-Time Scalable Diffusion Models (VFScale), a novel framework designed to solve complex reasoning problems. VFScale consists of innovations in both training objectives and inference scaling as shown in Fig. 2 to respectively address the two aforementioned challenges. (1) On the training side, VFScale incorporates a novel Linear-Regression Negative Contrastive Learning (LRNCL) objective and a KL regularization to improve the energy landscape, especially to better align the energy function with sample quality with LRNCL loss. The LRNCL objective samples two negatives [^1] Here, positive samples refer to data points, while negative samples correspond to noise-perturbed versions of the data. for each positive sample and fits a linear regression between their energy values and their L2 distances to the positive. It encourages a positively sloped linear relationship, aligning the energy landscape with distance to the ground truth. We refer to this alignment as performance-energy consistency. [^2] Detailed description for performance-energy consistency can be found in Appendix A. The KL regularization term backpropagates through the denoising process, smoothing the energy landscape. (2) On the inference side, we integrate the denoising process with a novel hybrid Monte Carlo Tree Search (hMCTS). It sequentially performs BoN and MCTS as the denoising proceeds. In the initial steps, BoN carries out parallel denoising without selecting to avoid premature exclusion of potential solutions, while the later MCTS denoising allows better navigation in the search space. Taken together, the hybrid MCTS denoising allows the model to significantly improve inference performance by scaling up the inference computational budget.
@@
We demonstrate the effectiveness of our VFScale’s training objective and scalable inference method on challenging reasoning tasks of Maze and Sudoku. In Sudoku tasks, our VFScale solves $43\%$ problems when conditioned on much less given digits (out-of-distribution condition), while the standard diffusion model only solves $30\%$ problems. In Maze tasks, trained with Maze sizes of up to $6\times 6$, VFScale successfully solves $88\%$ of substantially larger $15\times 15$ instances (see Fig. 1), whereas the standard diffusion model fails entirely.
@@
In summary, our contributions are as follows: (1) We introduce Verifier-free Test-time Scalable Diffusion Models (VFScale), a novel framework that can scale up test-time computation for better reasoning capability. (2) We introduce Linear-Regression Negative Contrastive Learning (LRNCL) and KL regularization into the training objective to improve the energy landscape. (3) We integrate a hybrid Monte Carlo Tree Search into the denoising process, enabling test-time scaling with better search efficiency. (4) We demonstrate the effectiveness of our method on challenging Sudoku and Maze datasets where the conditions are out-of-distribution or the problem sizes are much larger.
@@
Figure 2: Overview of VFScale. This figure illustrates the key aspects of VFScale by contrasting its training and inference strategies with those of the previous method. (1) To qualify the intrinsic energy of diffusion models as a verifier, VFScale introduces ${\mathcal{L}}_{\text{LRNCL}}$ and ${\mathcal{L}}_{\text{KL}}$ to improve the energy landscape during training. (2) In order for a higher search efficiency, VFScale proposes hybrid Monte Carlo Tree Search (hMCTS) that achieves a balance between best-of-$N$ and MCTS.
@@ 2 Related Work
Diffusion for Reasoning: The iterative refinement process inherent in diffusion models lends itself well to tasks requiring multi-step reasoning. du2022learning firstly formulated multi-step reasoning as an energy minimization problem. To approach harder problems with more complex energy landscapes, more compute budget is used for optimization steps. du2024learning formulated the energy-based model as a diffusion model, and the performance on harder tasks was significantly improved. While effective, the performance gain by scaling the compute of these methods quickly saturates as the number of sampling steps increases. Building upon du2024learning , VFScale instead explores number-of-sample scaling and proposes corresponding training and inference strategies.
Test-time Scaling of Diffusion: Improving diffusion model samples by increasing compute resources at inference time has emerged as an increasingly important direction. Early approaches focused on increasing the number of sampling steps, which leads to diminishing returns ho2020denoising ; karras2022elucidating . More recent work ma2025inferencetimescalingdiffusionmodels ; yoon2025montecarlotreediffusion has explored alternative scaling dimensions, such as increasing the number of sampled candidates, though these methods typically rely on external verifiers to guide solution selection. VFScale advances beyond these methods by leveraging the intrinsic energy function to guide test-time search, a capability further cultivated through customized training and inference strategies. These innovations make VFScale a versatile and scalable solution for real-world reasoning tasks.
@@ 3 Preliminary
The Denoising Diffusion Probabilistic Model (DDPM) ho2020denoising contains a predefined forward process to corrupt data into Gaussian noises, and a learnable reverse process to generate new samples from them. The forward process follows a Gaussian transition kernel $q(\bm{x}_{t}|\bm{x}_{t-1})=\mathcal{N}(\bm{x}_{t}|\sqrt{\alpha_{t}}\bm{x}_{t-1},(1-\alpha_{t})\bm{I}),t=1,\cdots,T$, where the noise schedule $\{\alpha_{t}\}_{t=1}^{T}$ and $T$ is set to make $\bm{x}_{T}$ follow approximately a standard Gaussian distribution. The reverse process can be learned to predict the noise from the corrupted data by minimizing
@@
where the expectation is w.r.t. $\bm{x}_{0}\sim p(\bm{x})$, $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$, and $t\sim\{1,...T\}$. $\bar{\alpha}_{t}:=\prod_{i=1}^{t}\alpha_{i}$. Eq. 1 is equivalent to optimizing a reweighted variational bound on negative log-likelihood. Without loss of generality, in this work, we parameterize $\bm{\epsilon}_{\theta}$ as the negative gradient of the energy function $\nabla_{\bm{x}_{t}}E_{\theta}(\bm{x}_{t},t)$ as in du2023reduce . To improve the energy landscape, du2024learning introduced a contrastive loss
@@
The reverse process starts with $\bm{x}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$, and iteratively applies the learned denoising net $\bm{\epsilon}_{\theta}$, where song2020denoising introduces an adjustable noise scale $\sigma_{t}$:
$\bm{x}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\frac{\bm{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{\theta}\left(\bm{x}_{t},t\right)}{\sqrt{\bar{\alpha}_{t}}}+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\bm{\epsilon}_{\theta}\left(\bm{x}_{t},t\right)+\sigma_{t}\bm{\epsilon}_{t},\,\,\bm{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (3)
Importantly, song2020denoising highlights that a diffusion model trained with $\{\alpha_{t}\}_{t=1}^{T}$ can be used to sample with $\{\alpha_{\tau_{s}}\}_{s=1}^{S}$, where $\tau$ is any increasing sub-sequence of $1,2,\cdots,T$ of length $S$. This variable spacing significantly accelerates sampling and will serve as an efficient approach for MCTS simulation in Section 4.3.
@@ 4.1 Problem setup
Let $\mathcal{D}=\{\mathcal{C},\mathcal{X}\}$ denote a dataset for a reasoning task, where each instance consists of an input ${\bm{c}}\in\mathbb{R}^{O}$ and its corresponding solution ${\bm{x}}\in\mathbb{R}^{M}$. Our goal is to model the complex reasoning problem as an optimization task to optimize an objective $\mathcal{J}({\bm{c}},{\bm{x}})$ and to capitalize on the diffusion model’s ability to iteratively refine the solution ${\bm{x}}$.
In this work, we leverage an energy-based diffusion model to learn an energy function $E_{\theta}({\bm{c}},{\bm{x}}_{t},t)$ [^3] For simplicity, we omit ${\bm{c}}$ in subsequent notations. to model the optimization objective $\mathcal{J}({\bm{c}},{\bm{x}})$. During inference, we apply test-time scaling by varying the number of samples and use the learned energy function as an intrinsic verifier, thereby eliminating the need for an external verifier. However, the reasoning performance gain by test-time scaling is highly dependent on both the quality of the energy landscape and the inference-time search algorithm. To address these challenges, we propose VFScale as shown in Fig. 2, with our solutions detailed in Section 4.2 and Section 4.3.
4.2 Training of VFScale
@@
Although the denoising loss ${\mathcal{L}}_{\text{MSE}}$ (Eq. 1) and the contrastive loss ${\mathcal{L}}_{\text{Contrast}}$ (Eq. 2) are effective, we find that simply scaling up inference budget through BoN processes yields marginal gain (Sec. 5.1). Through deeper analysis, we find that the core reason is the low quality of the energy landscape, espacially the lack of performance-energy consistency, which we address by introducing a novel Linear-Regression Negative Contrastive Learning loss and incorporating a KL regularization, which we detail as follows.
Linear-Regression Negative Contrastive Learning. Specifically, while the contrastive loss drives positive examples to the global energy minimum, it imposes no constraints on the relative energy ordering among negative samples. For example, it is likely that a negative example ${\bm{x}}_{t}^{--}$ that is further apart from the positive example ${\bm{x}}_{t}$ has lower energy than a negative example ${\bm{x}}_{t}^{-}$ that is nearer. Thus, the inference process can be stuck around ${\bm{x}}_{t}^{--}$ and can hardly move out. This will result in reduced performance-energy consistency, where lower energy at step $t$ does not necessarily correspond to a more accurate predicted solution $\hat{{\bm{x}}}_{0}$, as will be shown in Sec. 5.1.
This problem can not be easily remedied by increasing inference budget through BoN during test; instead, we explore a more fundamental way to solve this problem to shape the energy landscape by regularizing the energy among negative examples during training. The approach seeks to enforce consistency between the relative energy levels of negative samples and their corresponding performance quality, comprising the following key steps:
(1) Generate negative samples: Specifically, from a positive example ${\bm{x}}_{0}\sim p({\bm{x}})$, we sample two negative examples ${\bm{x}}_{0}^{-}$ and ${\bm{x}}_{0}^{--}$, [^4] Details to generate negative samples are in Appendix A. the latter has a larger L2 distance to ${\bm{x}}_{0}$. Specifically, the distance of ${\bm{x}}_{0}$, ${\bm{x}}_{0}^{-}$ and ${\bm{x}}_{0}^{--}$ to ${\bm{x}}_{0}$ are 00, $l_{2,0}^{-}=||{\bm{x}}_{0}^{-}-{\bm{x}}_{0}||_{2}^{2}$, and $l_{2,0}^{--}=||{\bm{x}}_{0}^{--}-{\bm{x}}_{0}||_{2}^{2}$, respectively.
(2) Obtain the energy for each sample: Then we obtain their corresponding noisy examples at step $t$ via ${\bm{x}}_{t}=\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}$, ${\bm{x}}_{t}^{-}=\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}^{-}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}$, and ${\bm{x}}_{t}^{--}=\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}^{--}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}$. Their energy are $E^{+}_{t}=E_{\theta}({\bm{x}}_{t},t)$, $E^{-}_{t}=E_{\theta}({\bm{x}}_{t}^{-},t)$, and $E^{--}_{t}=E_{\theta}({\bm{x}}_{t}^{--},t)$, respectively.
(3) Apply linear regression to get the slope and intercept: To encourage a landscape where the energy difference between negative samples and the positive sample becomes smaller as they approach the positive sample, we directly constrain the relationship between the energy of different negative samples and the positive sample using linear regression. Specifically, we first apply the linear regression algorithm [^5] The details of linear regression algorithm can be found in Appendix A. to fit linear function through the three points $\{(0,E_{t}^{+}),(l_{2,0}^{-},E_{t}^{-}),(l_{2,0}^{--},E_{t}^{--})\}$, which is characterized by the slope $k_{t}$ and the intercept $b_{t}$.
(4) Calculate the LRNCL loss: Then, we obtain the corresponding fitted points $\{(0,\hat{E}_{t}^{+}),(l_{2,0}^{-},\hat{E}_{t}^{-}),(l_{2,0}^{--},\hat{E}_{t}^{--})\}$ from the linear function. We then compute the Linear-Regression Negative Contrastive Learning (LRNCL) loss as:
${\mathcal{L}}_{\text{LRNCL}}={\mathbb{E}}_{{\bm{x}}_{0},{\bm{x}}_{0}^{-},{\bm{x}}_{0}^{--},{\bm{\epsilon}},t}\big{[}\max(0,\gamma-k_{t})\,+||E_{t}^{+}-\hat{E}_{t}^{+}||_{2}^{2}+||E_{t}^{-}-\hat{E}_{t}^{-}||_{2}^{2}+||E_{t}^{--}-\hat{E}_{t}^{--}||_{2}^{2}\big{]},$ (4)